OpenStudy (mythorius):

If the sinA = 3/5 and the cosA = 4/5, then what is tan A? 3/4 1 4/3 @zepdrix @KamiBug

OpenStudy (bobo-i-bo):

\[tanA=\frac{ sinA }{ cosA }\]

OpenStudy (bobo-i-bo):

By definition of tan

OpenStudy (kamibug):

Tangent = opp/adj Sine = opp/hyp Cosine = adj/ hyp Using the given values for sine and cosine, find opposite and hypotenuse to form tangent. :)

OpenStudy (mythorius):

Thank you both :3